A moving object slows down its motion from 12m/s to rest at a distance of 20m. Find the acceleration of the object(assumed constant)inside the barrel of the rifle. 




1
Expert's answer
2022-02-04T07:41:13-0500
"a=v^2/(2d)=12^2/(2·20)=3.6\text{ m/s}^2."